Grade upto college level General Physics

A stone is dropped from a height of 300m and at the same another stone is projected vertically upwards with a velocity of 100ms-1.Find when and where the two stones meet?

aditya kashyap
12 Years agoGrade upto college level

2 Answers

Yashwant
9 Years ago
Given the height with Wich the ball is left freely is 300m Let the two bodies will meet at a distance `x` meters from the heighest pont , let they meet in time `t` seconds X=ur+1/2gt^2X=5t^2
Sheharyaar
9 Years ago
for this question we can use relative speed concept
total distance d = 300m
relative speed = (100 × 0) m/s  [since the objects are moving in the opp direction]
therefore time t = d/s = 300/100 = 3s
so by  s = ut + ½ gt^2 = 0(3) + ½ * 10 * 3^2 = 5 * 9 =45 m
 
I hope you like my answer☺